Grade 11Mechanics

The object is thrown from the top of a building 49m high with initial velocity of 29.4 m/s making an angle of 30° with the horizontal. Determine the time taken by the object to hit the horizontal ground ,the distance from the foot of the building , also find the velocity with which it hits the ground

Yash Takodara
7 Years agoGrade 11

1 Answer

Srishti
7 Years ago
Time taken could be calculated by the formula …
T=2usin£/g...=2×29.4×sin30/10=2.9 seconds
 
Distance could be calculated by the range which is R=u^2sin2£/g
= (29.4)^2×2sin30cos30/10.=  69.14m